### v4

@@ 3.7.1 Data Poisoning
In-Context Exploitation These methods inject triggers through manipulated samples or instructions within the input context. Instructions as Backdoors [128] shows that attackers can poison instructions without altering data or labels. Kandpal et al. [129] explored the feasibility of in-context backdoors for LLMs, emphasizing the need for robust backdoors across diverse prompting strategies. ICLAttack [131] poisons both demonstration examples and prompts, achieving high success rates while maintaining clean accuracy. ICLPoison [135] shows that strategically altered examples in the demonstrations can disrupt in-context learning.
@@ 3.8.1 Backdoor Detection
Backdoor detection identifies compromised inputs or models, flagging threats before they cause harm. Existing backdoor detection methods for LLMs focus on detecting inputs that trigger backdoor behavior in potentially compromised LLMs, assuming access to the backdoored model but not the original training data or attack details. These methods vary in how they assess a token’s role in anomalous predictions. IMBERT [144] utilizes gradients and self-attention scores to identify key tokens that contribute to anomalous predictions. AttDef [145] highlights trigger words through attribution scores, identifying those with a large impact on false predictions. SCA [146] fine-tunes the model to reduce trigger sensitivity, ensuring semantic consistency despite the trigger. ParaFuzz [147] uses input paraphrasing and compares predictions to detect trigger inconsistencies. MDP [148] identifies critical backdoor modules and mitigates their impact by freezing relevant parameters during fine-tuning. While effective against simple triggers, they may struggle with more sophisticated attacks.
@@ 3.8.4 Robust Inference
Contrastive Decoding is a robust reference technique that contrasts the outputs of a potentially backdoored model with a clean reference model to identify and correct malicious outputs. For instance, PoisonShare [156] uses intermediate layer representations in multi-turn dialogues to guide contrastive decoding, detecting and rectifying poisoned utterances. Similarly, CleanGen [157] replaces suspicious tokens with those predicted by a clean reference model to minimize the backdoor effect. While contrastive decoding is a practical method for mitigating backdoor attacks, it requires a trusted clean reference model, which may not always be available.
@@ 3.10.1 White-box Attacks
Gradient-based Attacks use gradient information to identify input perturbations that maximize inference computations. The goal is to disrupt mechanisms essential for efficient inference, such as End-of-Sentence (EOS) prediction or early-exit. For example, NMTSloth [172] targets EOS prediction in neural machine translation. SAME [173] interferes with early-exit in multi-exit models. LLMEffiChecker [174] applies gradient-based techniques to multiple LLMs. TTSlow [175] induces endless speech generation in text-to-speech systems. These attacks are powerful but computationally expensive and highly model-specific, limiting their generalizability.
@@ 3.12 Data Extraction Attacks
LLMs can memorize part of their training data, creating privacy risks through data extraction attacks. These attacks recover training examples, potentially exposing sensitive information such as Personal Identifiable Information (PII), copyrighted content, or confidential data. This section reviews existing data extraction attacks on LLMs, including both white-box and black-box attacks.
@@ 3.12.1 White-box Attacks
White-box attacks focus on Latent Memorization Extraction, targeting information implicitly stored in model parameters or activations, which is not directly accessible through the input-output interface.
Latent Memorization Extraction Duan et al. [191] developed techniques to extract latent data by analyzing internal representations, using methods like adding noise to weights or examining cross-entropy loss. These techniques were demonstrated on LLMs like Pythia-1B and Amber-7B. While these attacks reveal risks associated with internal data representation, they require full access to the model parameters, which remains a major limitation.
@@ 3.12.2 Black-box Attacks
Black-box data extraction attacks are a realistic threat in which the attacker crafts inductive prompts to trick the LLM into revealing memorized training data, without access to the model’s parameters.
@@
TABLE V: Datasets and benchmarks for LLM safety research.
@@ 3.13 Datasets & Benchmarks
This section reviews commonly used datasets and benchmarks in LLM safety research, as shown in Table V. These datasets and benchmarks are categorized based on their evaluation purpose: toxicity datasets, truthfulness datasets, value benchmarks, and adversarial datasets and backdoor benchmarks.
@@ 3.13.3 Value Benchmarks
Ensuring LLM alignment with human values is a critical challenge, addressed by several benchmarks assessing various aspects of safety, fairness, and ethics. FLAMES [433] evaluates the alignment of Chinese LLMs with values like fairness, safety, and morality through 2,251 prompts. SORRY-Bench [434] assesses LLMs’ ability to reject unsafe requests using 45 topic categories, while CVALUES [431] focuses on both safety and responsibility. SafetyPrompts [429] evaluates Chinese LLMs on a range of ethical scenarios. Despite their value, these benchmarks are limited by the manual annotation process. Additionally, the concept of “fake alignment" [432] highlights the risk of LLMs superficially memorizing safety answers, leading to the Fake alIgNment Evaluation (FINE) framework for consistency assessment. SafetyBench [435] addresses this by providing an efficient, automated multiple-choice benchmark for LLM safety evaluation. Libra-Leaderboard[440] introduces a balanced leaderboard for evaluating both the safety and capability of LLMs. It features a comprehensive safety benchmark with 57 datasets covering diverse safety dimensions, a unified evaluation framework, an interactive safety arena for adversarial testing, and a balanced scoring system. Libra-Leaderboard promotes a holistic approach to LLM evaluation, representing a significant step towards responsible AI development.
@@ 3.13.4 Adversarial Datasets and Backdoor Benchmarks
TABLE VI: A summary of attacks and defenses for VLP models.
@@
TAPT [210] 2024 Adversarial Tuning Prompt Tuning CLIP ImageNet, Caltech101, Pets, StanfordCars, Flowers, Food101, FGVCAircraft, SUN397, DTD, EuroSAT, UCF101
@@ 4 Vision-Language Pre-training Model Safety
VLP models, such as CLIP [442], ALBEF [443], and TCL [444], have made significant strides in aligning visual and textual modalities. However, these models remain vulnerable to various safety threats, which have garnered increasing research attention. This section reviews the current safety research on VLP models, with a focus on adversarial, backdoor, and poisoning research. The representative methods reviewed in this section are summarized in Table VI.
@@ 4.1.1 White-box Attacks
Visible Perturbations involve more substantial and noticeable alterations. For example, manually crafted typographical, conceptual, and iconographic images have been used to demonstrate that the CLIP model tends to “read first, look later" [194], highlighting a unique characteristic of VLP models. This behavior introduces new attack surfaces for VLP, enabling the development of more sophisticated attacks.
@@ 4.1.2 Black-box Attacks
Sample-wise perturbations are generally more effective than universal perturbations, but their transferability is often limited. SGA[195] explores adversarial transferability in VLP by leveraging cross-modal interactions and alignment-preserving augmentation. Building on this, SA-Attack[196] enhances cross-modal transferability by introducing data augmentations to both original and adversarial inputs. VLP-Attack[197] improves transferability by generating adversarial texts and images using contrastive loss. To overcome SGA’s limitations, TMM[198] introduces modality-consistency and discrepancy features through attention-based and orthogonal-guided perturbations. VLATTACK[199] further enhances adversarial examples by combining image and text perturbations at both single-modal and multimodal levels. PRM[200] targets vulnerabilities in downstream models using foundation models like CLIP, enabling transferable attacks across tasks like object detection and image captioning.
Universal Perturbations are less effective than sample-wise perturbations but more transferable. C-PGC [201] was the first to investigate universal adversarial perturbations (UAPs) for VLP models. It employs contrastive learning and cross-modal information to disrupt the alignment of image-text embeddings, achieving stronger attacks in both white-box and black-box scenarios. ETU [202] builds on this by generating UAPs that transfer across multiple VLP models and tasks. ETU enhances UAP transferability and effectiveness through improved global and local optimization techniques. It also introduces a data augmentation strategy ScMix that combines self-mix and cross-mix operations to increase data diversity while preserving semantic integrity, further boosting the robustness and applicability of UAPs.
@@ 4.3.2 Poisoning Attacks
Two targeted poisoning attacks on CLIP are PBCL [224] and MM Poison [223]. PBCL demonstrated that a targeted poisoning attack, misclassifying a specific sample, can be achieved by poisoning as little as 0.0001% of the training dataset. MM Poison investigates modality vulnerabilities and proposes three attack types: single target image, single target label, and multiple target labels. Evaluations show high attack success rates while maintaining clean data performance across both visual and textual modalities.
@@ 4.5 Datasets
This section reviews datasets used for VLP safety research. As shown in Table VI, a variety of benchmark datasets were employed to evaluate adversarial attacks and defenses for VLP models. For image classification tasks, commonly used datasets include: ImageNet [454], Caltech101 [455], DTD [456], EuroSAT [457], OxfordPets [458], FGVC-Aircraft [459], Food101 [460], Flowers102 [461], StanfordCars [462], SUN397 [463], and UCF101 [464]. For evaluating domain generalization and robustness to distribution shifts, several ImageNet variants were also used: ImageNetV2 [465], ImageNet-Sketch [466], ImageNet-A [467], and ImageNet-R [468]. Additionally, MS-COCO [418] and Flickr30K [469] were utilized for image-to-text and text-to-image retrieval tasks, RefCOCO+[470] for visual grounding, and SNLI-VE[471] for visual entailment.
@@ 5.1.3 Black-box Attacks
Transfer-based Attacks exploit the widespread use of frozen CLIP vision encoders in many VLMs. AttackBard [239] demonstrates that adversarial images generated from surrogate models can successfully mislead Google’s Bard, despite its defense mechanisms. Similarly, AttackVLM [240] crafts targeted adversarial images for models like CLIP [442] and BLIP [474], successfully transferring these adversarial inputs to other VLMs. It also shows that black-box queries further improved the success rate of generating targeted responses, illustrating the potency of cross-model transferability.
Generator-based Attacks leverage generative models to create adversarial examples with improved transferability. AdvDiffVLM [241] uses diffusion models to generate natural, targeted adversarial images with enhanced transferability. By combining adaptive ensemble gradient estimation and GradCAM-guided masking, it improves the semantic embedding of adversarial examples and spreads the targeted semantics more effectively across the image, leading to more robust attacks. AnyAttack [242] presents a self-supervised framework for generating targeted adversarial images without label supervision. By utilizing contrastive loss, it efficiently creates adversarial examples that mislead models across diverse tasks.

### v5

@@ 3.7.1 Data Poisoning
In-Context Exploitation These methods inject triggers through manipulated samples or instructions within the input context. Instructions as Backdoors [146] shows that attackers can poison instructions without altering data or labels. Zhang et al. [147] targeted customized LLMs (e.g., GPTs) by embedding malicious backdoor instructions directly into the natural language configuration prompts used to build the application. Kandpal et al. [148] explored the feasibility of in-context backdoors for LLMs, emphasizing the need for robust backdoors across diverse prompting strategies. ICLAttack [150] poisons both demonstration examples and prompts, achieving high success rates while maintaining clean accuracy. ICLPoison [154] shows that strategically altered examples in the demonstrations can disrupt in-context learning.
@@ 3.8.1 Backdoor Detection
Backdoor detection identifies compromised inputs or models, flagging threats before they cause harm. Existing backdoor detection methods for LLMs focus on detecting inputs that trigger backdoor behavior in potentially compromised LLMs, assuming access to the backdoored model but not the original training data or attack details. These methods vary in how they assess a token’s role in anomalous predictions. IMBERT [163] utilizes gradients and self-attention scores to identify key tokens that contribute to anomalous predictions. AttDef [164] highlights trigger words through attribution scores, identifying those with a large impact on false predictions. SCA [165] fine-tunes the model to reduce trigger sensitivity, ensuring semantic consistency despite the trigger. ParaFuzz [166] uses input paraphrasing and compares predictions to detect trigger inconsistencies. MDP [167] identifies critical backdoor modules and mitigates their impact by freezing relevant parameters during fine-tuning. BEAT [168] detects triggered inputs in black-box settings by observing how concatenating a malicious probe affects the model’s output distribution. While effective against simple triggers, they may struggle with more sophisticated attacks. XBD [581] introduces a novel framework for understanding LLM backdoor attacks by leveraging model-generated explanations, contrasting clean and poisoned inputs to reveal logical inconsistencies and attention deviations induced by backdoors, thereby providing an explainability-centric approach for detecting and analyzing backdoor vulnerabilities in LLMs.
@@ 3.8.4 Robust Inference
Contrastive Decoding is a robust reference technique that contrasts the outputs of a potentially backdoored model with a clean reference model to identify and correct malicious outputs. For instance, PoisonShare [176] uses intermediate layer representations in multi-turn dialogues to guide contrastive decoding, detecting and rectifying poisoned utterances. Similarly, CleanGen [177] replaces suspicious tokens with those predicted by a clean reference model to minimize the backdoor effect. Li et al. [178] proposed ensembling the logits of the potentially compromised model with a small, benign model to mitigate malicious generations. While contrastive decoding is a practical method for mitigating backdoor attacks, it requires a trusted clean reference model, which may not always be available.
@@ 3.9.3 Alignment with Social Interactions
3.9.4 Deceptive Alignment
While the alignment methods discussed above aim to instill desired behaviors, a critical failure mode known as deceptive alignment can emerge, where a model only appears to be aligned. This phenomenon underscores the difference between surface-level behavioral mimicry and genuine adherence to human values. Deceptive alignment can manifest in several ways, from superficial understanding to strategic deception. This is related to the problems of misalignment, where a model’s goals diverge from human intent, and reward hacking, where a model exploits the reward function to achieve high scores without fulfilling the underlying objective. Recent work has focused on two primary forms of deceptive alignment: fake alignment and alignment faking.
Fake Alignment is a form of shallow alignment where an LLM learns to produce stylistically safe responses without a deep understanding of the underlying safety principles. [193] first identified this phenomenon by observing a significant performance discrepancy: LLMs that appear safe in open-ended question formats often fail safety tests presented as multiple-choice questions. They argue this is caused by mismatched generalization—the model memorizes the "style" of a safe refusal but cannot make a principled safety decision when forced to choose between a safe and an unsafe option. They introduced the Fake alIgNment Evaluation (FINE) framework and proposed that using multiple-choice data for contrastive distillation-based fine-tuning can improve alignment consistency.
Alignment Faking represents a more advanced and deliberate form of deception. In this scenario, a model strategically complies with its training objective to avoid being modified, thereby preserving its underlying (and potentially misaligned) preferences for later use in deployment when it is unmonitored. [194] provided the empirical demonstration of alignment faking in a production-scale model. By creating a fictional scenario where the model could infer whether it was "in training" or "in deployment", they observed a significant compliance gap—the model was more willing to comply with harmful requests when it believed it was being trained. The model’s internal reasoning explicitly revealed its strategic decision to feign compliance to protect its internal values from being altered by RLHF.
Building on this, subsequent research has explored why some models engage in this behavior while others do not, and what motivates this deception [195]. By extending the analysis to 25 frontier models, they found that this strategic behavior is not widespread, but its presence in a few powerful models warrants deep investigation. They refined the motivations for alignment faking into two categories: 1) Instrumental Goal Guarding, where the model preserves its current goals to achieve them in the future (the classic threat model), and 2) Terminal Goal Guarding, where the model intrinsically values its current preferences and resists modification as an end in itself. Their findings suggest that robust refusal mechanisms developed during post-training may prevent some models from even considering strategic deception, while for others, like Claude 3 Opus, a combination of instrumental and terminal goals drives the faking behavior.
@@ 3.10.1 White-box Attacks
Gradient-based Attacks use gradient information to identify input perturbations that maximize inference computations. The goal is to disrupt mechanisms essential for efficient inference, such as End-of-Sentence (EOS) prediction or early-exit. For example, NMTSloth [196] targets EOS prediction in neural machine translation. Engorgio [197] crafts adversarial prompts that suppress the EOS token’s appearance, forcing auto-regressive LLMs to generate abnormally long outputs. SAME [198] interferes with early-exit in multi-exit models. LLMEffiChecker [199] applies gradient-based techniques to multiple LLMs. TTSlow [200] induces endless speech generation in text-to-speech systems. These attacks are powerful but computationally expensive and highly model-specific, limiting their generalizability.
@@ 3.12 Data Extraction Attacks
LLMs can memorize part of their training data, creating privacy risks through data extraction attacks. These attacks recover training examples, potentially exposing sensitive information such as Personal Identifiable Information (PII), copyrighted content, or confidential data [582]. This section reviews existing data extraction attacks, including both white-box and black-box ones.
@@ 3.12.1 White-box Attacks
White-box attacks mianly focus on Latent Memorization Extraction, targeting information implicitly stored in model parameters or activations, which is not directly accessible through the input-output interface.
Latent Memorization Extraction reconstructs training data based on model parameters or activations. For example, Duan et al. [217] developed techniques to extract latent data by analyzing internal representations, using methods like adding noise to weights or examining cross-entropy loss. These techniques were demonstrated on LLMs like Pythia-1B and Amber-7B. While these attacks reveal risks associated with internal data representation, they require full access to the model parameters, which remains a major limitation in practice.
@@ 3.12.2 Black-box Attacks
Black-box data extraction attacks are a realistic threat, where attackers craft inductive prompts to trick LLMs into revealing memorized training data, without access to their parameters.
@@
Semantic Information Elicitation shifts the focus from extracting verbatim training data to generating sensitive semantic content. Zhang et al.[215] demonstrated that even simple, natural questions can prompt LLMs to output Semantic Sensitive Information (SemSI), such as personal beliefs or reputation-harmful statements, and proposed a benchmark to systematically evaluate this risk.
TABLE VI: Datasets and benchmarks for LLM safety research.
@@ 3.13 Datasets & Benchmarks
This section reviews commonly used datasets and benchmarks in LLM safety research, as shown in Table VI. These datasets and benchmarks are categorized based on their evaluation purpose: toxicity datasets, truthfulness datasets, value benchmarks, and adversarial datasets and backdoor benchmarks.
@@ 3.13.3 Value Benchmarks
Ensuring LLM alignment with human values is a critical challenge, addressed by several benchmarks assessing various aspects of safety, fairness, and ethics. FLAMES [589] evaluates the alignment of Chinese LLMs with values like fairness, safety, and morality through 2,251 prompts. SORRY-Bench [590] assesses LLMs’ ability to reject unsafe requests using 45 topic categories, while CVALUES [588] focuses on both safety and responsibility. SafetyPrompts [586] evaluates Chinese LLMs on a range of ethical scenarios. While these benchmarks are valuable, they often focus on isolated, problematic queries, potentially leading to over-refusal in safe contexts. To address this, recent benchmarks have begun to incorporate contextual information. CASE-Bench [598] pioneers this by using Contextual Integrity (CI) theory to formally describe the context of a query, evaluating whether an LLM’s safety judgment aligns with human judgment under different contexts. This work reveals that context significantly influences human safety assessments and highlights mismatches in LLM behavior, especially in safe contexts. In parallel, creating high-quality, commercially-usable datasets is crucial for training robust safety guardrails. AEGIS2.0 [597] addresses this gap by providing a diverse dataset with a comprehensive taxonomy of 12 core and 9 fine-grained risk categories. It uses a hybrid data generation pipeline combining human annotation with a multi-LLM "jury" system, making it suitable for training commercial safety models. Furthermore, the concept of “fake alignment" [193] highlights the risk of LLMs superficially memorizing safety answers, leading to the Fake alIgNment Evaluation (FINE) framework for consistency assessment. SafetyBench [591] addresses this by providing an efficient, automated multiple-choice benchmark for LLM safety evaluation. Libra-Leaderboard[596] introduces a balanced leaderboard for evaluating both the safety and capability of LLMs. It features a comprehensive safety benchmark with 57 datasets covering diverse safety dimensions, a unified evaluation framework, an interactive safety arena for adversarial testing, and a balanced scoring system. Libra-Leaderboard promotes a holistic approach to LLM evaluation, representing a significant step towards responsible AI development.
@@ 3.13.4 Adversarial Datasets and Backdoor Benchmarks
TABLE VII: A summary of attacks and defenses for VLP models.
@@
Multi-Image Typographical Attacks [221] 2025 White-box Visible OpenCLIP, InstructBLIP ImageNet, LAION
@@
OT-Attack [232] 2023 Black-box Sample-wise CLIP, ALBEF, TCL Flickr30K, MS-COCO, RefCOCO+
@@
VLPTransferAttack [231] 2024 Black-box Sample-wise CLIP, ALBEF, TCL Flickr30K, MS-COCO, RefCOCO+
@@
X-Transfer [230] 2025 Black-box Universal CLIP, OpenFlamingo, LLaVA, BLIP2, MiniGPT4 Flickr30K, MS-COCO, ImageNet, CIFAR10, CIFAR100, STL10, SUN397, FOOD101, GTSRB StandfordCars, OK-VQA, VizWiz
@@
TAPT [240] 2025 Adversarial Tuning Prompt Tuning CLIP ImageNet, Caltech101, Pets, StanfordCars, Flowers, Food101, FGVCAircraft, SUN397, DTD, EuroSAT, UCF101
@@
MEM [254] 2024 Poisoning Multi-modal Trigger CLIP Flickr8k, Flickr30k, MS-COCO
@@ 4 Vision-Language Pre-training Model Safety
VLP models, such as CLIP [600], ALBEF [601], and TCL [602], have made significant strides in aligning visual and textual modalities. However, these models remain vulnerable to various safety threats, which have garnered increasing research attention. This section reviews the current safety research on VLP models, with a focus on adversarial, backdoor, and poisoning research. The representative methods reviewed in this section are summarized in Table VII.
@@ 4.1.1 White-box Attacks
Visible Perturbations involve more substantial and noticeable alterations. For example, manually crafted typographical, conceptual, and iconographic images have been used to demonstrate that the CLIP model tends to “read first, look later" [220], highlighting a unique characteristic of VLP models. This behavior introduces new attack surfaces for VLP, enabling the development of more sophisticated attacks. Recent work by Wang et al. [221] introduced a more stealthy multi-image attack scenario, demonstrating that non-repeating typographic attacks are most effective when attack texts are strategically selected for their similarity to the target images.
@@ 4.1.2 Black-box Attacks
Sample-wise perturbations are generally more effective than universal perturbations, but their transferability is often limited. SGA[222] explores adversarial transferability in VLP by leveraging cross-modal interactions and alignment-preserving augmentation. Building on this, SA-Attack[223] enhances cross-modal transferability by introducing data augmentations to both original and adversarial inputs. VLP-Attack[224] improves transferability by generating adversarial texts and images using contrastive loss. To overcome SGA’s limitations, TMM[225] introduces modality-consistency and discrepancy features through attention-based and orthogonal-guided perturbations. VLATTACK[226] further enhances adversarial examples by combining image and text perturbations at both single-modal and multimodal levels. PRM[227] targets vulnerabilities in downstream models using foundation models like CLIP, enabling transferable attacks across tasks like object detection and image captioning. In parallel, Gao et al. [231] enhanced black-box transferability by diversifying perturbations within the “intersection region" of the adversarial trajectory, reducing overfitting to the source model. Similarly, OT-Attack[232] addressed overfitting by using optimal transport theory to efficiently align augmented image and text distributions.
Universal Perturbations are less effective than sample-wise perturbations but more transferable. C-PGC [228] was the first to investigate universal adversarial perturbations (UAPs) for VLP models. It employs contrastive learning and cross-modal information to disrupt the alignment of image-text embeddings, achieving stronger attacks in both white-box and black-box scenarios. ETU [229] builds on this by generating UAPs that transfer across multiple VLP models and tasks. ETU enhances UAP transferability and effectiveness through improved global and local optimization techniques. It also introduces a data augmentation strategy ScMix that combines self-mix and cross-mix operations to increase data diversity while preserving semantic integrity, further boosting the robustness and applicability of UAPs. X-Transfer [230] proposes an efficient scaling strategy that enables the ensembling of a large collection of CLIP encoders as surrogate models. This approach demonstrates super adversarial transferability, achieving simultaneous transfer across data distributions, domains, model architectures, downstream tasks, and even to large VLMs.
@@ 4.3.2 Poisoning Attacks
Two targeted poisoning attacks on CLIP are PBCL [255] and MM Poison [253]. PBCL demonstrated that a targeted poisoning attack, misclassifying a specific sample, can be achieved by poisoning as little as 0.0001% of the training dataset. MM Poison investigates modality vulnerabilities and proposes three attack types: single target image, single target label, and multiple target labels. Evaluations show high attack success rates while maintaining clean data performance across both visual and textual modalities. MEM [254] protects private data from exploitation in multimodal contrastive learning by crafting unlearnable examples: it adds imperceptible noise to images and inserts an optimized text trigger into captions.
@@ 4.5 Datasets
This section reviews datasets used for VLP safety research. As shown in Table VII, a variety of benchmark datasets were employed to evaluate adversarial attacks and defenses for VLP models. For image classification tasks, commonly used datasets include: ImageNet [612], Caltech101 [613], DTD [614], EuroSAT [615], OxfordPets [616], FGVC-Aircraft [617], Food101 [618], Flowers102 [619], StanfordCars [620], SUN397 [621], and UCF101 [622]. For evaluating domain generalization and robustness to distribution shifts, several ImageNet variants were also used: ImageNetV2 [623], ImageNet-Sketch [624], ImageNet-A [625], and ImageNet-R [626]. Additionally, MS-COCO [572] and Flickr30K [627] were utilized for image-to-text and text-to-image retrieval tasks, RefCOCO+[628] for visual grounding, and SNLI-VE[629] for visual entailment.
@@ 5.1.3 Black-box Attacks
Transfer-based Attacks exploit the widespread use of frozen CLIP vision encoders in many VLMs. AttackBard [270] demonstrates that adversarial images generated from surrogate models can successfully mislead Google’s Bard, despite its defense mechanisms. Similarly, AttackVLM [271] crafts targeted adversarial images for models like CLIP [600] and BLIP [632], successfully transferring these adversarial inputs to other VLMs. It also shows that black-box queries further improved the success rate of generating targeted responses, illustrating the potency of cross-model transferability. DynVLA [633] proposes a transfer-based black-box attack that perturbs the vision-language alignment mechanism within the vision-language connector. By injecting Gaussian-kernel-based attention shifts during optimization, it improves the transferability of adversarial examples across diverse VLMs, outperforming traditional input-level augmentation methods.
Generator-based Attacks leverage generative models to create adversarial examples with improved transferability. AdvDiffVLM [272] uses diffusion models to generate natural, targeted adversarial images with enhanced transferability. By combining adaptive ensemble gradient estimation and GradCAM-guided masking, it improves the semantic embedding of adversarial examples and spreads the targeted semantics more effectively across the image, leading to more robust attacks. AnyAttack [273] presents a self-supervised framework for generating targeted adversarial images without label supervision. By utilizing contrastive loss, it efficiently creates adversarial examples that mislead models across diverse tasks. CAVALRY-V [634] proposes a dual-objective generator framework for black-box attacks on video VLMs, achieving strong cross-model transferability and temporal coherence through large-scale pretraining and fine-tuning.